OpenStudy (anonymous):

pleasee help!!! ''identify the graph of the equation 2x^2+9xy+14y^2-5=0. then find θ''

OpenStudy (anonymous):

@precal do you know how to do this??

OpenStudy (precal):

you are looking for theta?

OpenStudy (precal):

this looks like a conic

OpenStudy (anonymous):

@jdoe0001 do u know how to do this?

OpenStudy (precal):

not sure why you need theta? What course is this for?

OpenStudy (anonymous):

precal

OpenStudy (anonymous):

the formula im suppose to use is tan2theta=b/a-c

OpenStudy (anonymous):

ive done this so far: tan2θ=9/2-14 tan2θ=-0.75 don't know what to do after that..:(

OpenStudy (precal):

you said this is an ellipse, correct

OpenStudy (anonymous):

yeaah

OpenStudy (precal):

\[\tan^2 \theta \]

OpenStudy (precal):

is that correct

OpenStudy (anonymous):

the answer is θ=-18 degrees

OpenStudy (precal):

hope that helps. I have never looked for the angle of rotation before so I am unable to help you sorry

OpenStudy (anonymous):

ok thanks anyway :)

OpenStudy (precal):

just sub your solution into your equation and you should be able to see if it works at all. Don't forget your even and odd identities